# A Bayesian framework for active object recognition, pose estimation and shape transfer learning through touch

### v1

@@ III-I Active Data Acquisition
As tactile data points at each time step only cover a fraction of the object surface, active data acquisition is required to address the uncertainty and ambiguity due to incomplete information. The process is divided into two steps: target point selection and contact enforcement. The idea is to first determine a target point, and start exploration from the candidate target point until a contact point is found.
III-I1 Target Point Selection
@@
One way to select the target point is to calculate the posterior variance for each vertex on the reconstructed surface based on equation (31), then the point with the highest posterior variance will be selected as the target point $\overline{\boldsymbol{x}}_{t+1}$ for the next time step, i.e.,
$\text{var}(\boldsymbol{x}^{*})=\boldsymbol{\Sigma}^{*}_{0,0}$ (39)
$\overline{\boldsymbol{x}}_{t+1}=\operatorname*{argmax}_{\boldsymbol{x}^{*}\in\boldsymbol{S}^{*}}\left\{\text{var}(\boldsymbol{x}^{*})\right\}$ (40)
$\boldsymbol{S}^{*}\subset\left\{\boldsymbol{x}^{*}\in R^{3}\;|\;\boldsymbol{\mu}^{*}(\boldsymbol{x}^{*})=0\right\},$ (41)
@@
with $\boldsymbol{S}^{*}$ a finite subset of the zero-level set of the posterior GPIS, which can be obtained through the marching cubes algorithm [33].
@@
However, this exploration procedure requires the framework to update the GPIS every time step, which is computationally expensive. To alleviate this issue, the GPIS will not be updated and the framework will use the MLE shape from PF as estimated shape when the object is considered to be a known object. An exploration based on the directed Hausdorff distance (DHD) is used to guide the exploration as variance from the GPIS is not available. The point of interest for the next time step $\overline{\boldsymbol{x}}_{t+1}$ can be selected on the MLE surface by finding the point that yields the largest DHD, i.e.,
$\overline{\boldsymbol{x}}_{t+1}=\operatorname*{argmax}_{\boldsymbol{x}_{m}\in\boldsymbol{M}}\min_{\boldsymbol{x}_{c}\in\boldsymbol{X}_{c}}\lVert\boldsymbol{x}_{m}-\boldsymbol{x}_{c}\rVert_{2},$ (42)
@@
where $\boldsymbol{M}$ represents the set of all vertices of the MLE shape, and $\boldsymbol{X}_{c}$ denotes the set of observed contact points. The point of interest obtained through equation (42) is the point on the MLE shape with the largest distance to its nearest data point. And DHD is defined as follows,
$\boldsymbol{d}_{H}(\boldsymbol{A},\boldsymbol{B})=\max_{\boldsymbol{a}\in\boldsymbol{A}}\min_{\boldsymbol{b}\in\boldsymbol{B}}\lVert\boldsymbol{a}-\boldsymbol{b}\rVert_{2}$ (43)
where $\boldsymbol{A}$ and $\boldsymbol{B}$ are two point sets.
@@
III-I2 Contact Enforcement
With equations (40) and (42), a candidate target point is obtained at each time step, but there is no guarantee that it would lead to contact. On the contrary, since the target point is near the furthest point on the estimated surface from the existing data point, it often fails to establish contact at the proposed target location. A simple solution to guarantee contact at each time step is to perform contour following, namely, the tactile sensor will move towards the target point locally on the tangent plane at each time step. However, this approach might take a long time to reach the target location and fail to capture the global context efficiently. As the selected target point is either on the GPIS or the MLE shape, confirming the occupancy near the target point can provide information that could greatly impact the belief of the framework. Negative information (non-contact points) around the target point implies the current estimation is likely to be wrong, whereas positive information (contact) provides evidence of the current estimation.
Figure 3: Example of the contact enforcement procedure. Starting from the target point (red dot), the sensor first moves towards the interior of the MLE shape (purple dot). If no contact is found, the sensor moves towards the closest existing contact point (green dot). If no contact is established either, the closest knwon contact point is contacted again. From there, the algorithm takes small steps on the surface along the local tangent plane, towards the target point while remaining in contact with the surface. After a short distance, a new contact point will ultimately be recorded (yellow dot).
Fig. 3 sketches the contact enforcement procedure; see caption for details. During the contact enforcement procedure, most waypoints only contain negative information. Naively taking all the negative information data into account will lead to the negative information dominating the likelihood function, while the contact points would have less impact. To compensate for the imbalanced data issue, only target points that are far from the final contact points will be registered as negative information. In this manner, the number of data points that contain negative information is no larger than the number of data points with positive information.
Alternatively, when no prior is available, a rapidly exploring random tree (RRT) based exploration procedure can be adopted. It simply chooses a random point in the workspace, and it expands in that direction from the closest existing contact point. This requires no prior knowledge or estimation of the global shape as the expansion is local.
III-J Termination criterion
@@
$\boldsymbol{d}_{H}(\boldsymbol{M},\boldsymbol{X}_{c})\leq\epsilon$ (44)
The algorithm terminates if $\boldsymbol{d}_{H}(\boldsymbol{M},\boldsymbol{X}_{c})$ is lower than a small value $\epsilon$, where $\boldsymbol{M}$ represents the set of all vertices of the MLE shape, and $\boldsymbol{X}_{c}$ denotes the set of observed contact points. Intuitively, if any point on the estimated surface is at most $\epsilon$ away from its closest existing contact point, the program terminates. In other words, to terminate the exploration, the contact points have to cover the estimated object surface with a certain density related to $\epsilon$. The threshold $\epsilon$ can be interpreted as a level of detail parameter: lower $\epsilon$ results in more data being collected, albeit with longer exploration time. Similarly, for novel objects, the termination criterion is
$\boldsymbol{d}_{H}(\boldsymbol{\boldsymbol{S}^{*}},\boldsymbol{X}_{c})\leq\epsilon,$ (45)
with $\boldsymbol{S}^{*}$ as the zero level set of the GPIS. Although the posterior variance of the GPIS can also be used for the same purpose, the proposed criterion $\boldsymbol{d}_{H}(\boldsymbol{\boldsymbol{S}^{*}},\boldsymbol{X}_{c})$ has the advantage of being independent of the choice of kernel parameters of the Gaussian process and the type of test objects at hand.
@@
Initialization: Initialize the particle filter with the first contact point $\boldsymbol{x}_{1}$ and tactile observation $\boldsymbol{d}_{1}$.
while $\boldsymbol{d}_{H}(\boldsymbol{\boldsymbol{S}^{*}},\boldsymbol{X}_{c})>\epsilon$ do
      Update belief using the particle filter, extract the MLE particle $(\boldsymbol{z}^{*},\overline{w}^{*})$ and its model evidence $p(\boldsymbol{D}|\boldsymbol{z}^{*},\boldsymbol{X})$
       if $p(\boldsymbol{D}|\boldsymbol{z}^{*},\boldsymbol{X})<(\frac{1}{Z}0.90)^{n_{pos}}*(0.50)^{n_{neg}}$ then
             Update GPIS and select the next target point $\overline{\boldsymbol{x}}_{t+1}$ from the zero level set of GPIS $\boldsymbol{S}^{*}$ using eq. (40)
@@
             Select the next target point $\overline{\boldsymbol{x}}_{t+1}$ from the set of MLE surface $\boldsymbol{M}$ using eq. (42)
@@
      Establish contact using the exploration procedure and add data points to $\boldsymbol{X}$ and data to $\boldsymbol{D}$.
       Update $\boldsymbol{d}_{H}(\boldsymbol{\boldsymbol{S}^{*}},\boldsymbol{X}_{c})$ based on eq. (43)
@@
III-K Shape similarity metric
@@
$\boldsymbol{d}_{TH}(\boldsymbol{A},\boldsymbol{B})=\max(\boldsymbol{d}_{H}(\boldsymbol{A},\boldsymbol{B}),\boldsymbol{d}_{H}(\boldsymbol{B},\boldsymbol{A}))$ (46)
This metric is used in the evaluation of the proposed framework. The pose estimation error is measured by the TWD between the MLE shape from the PF and the actual shape of the test object. Similarly, the shape reconstruction error is measured by the TWD distance between the reconstructed shape from the GPIS and the actual shape of the test object.
@@ IV-A Experiment setup
To show that the framework can address both known and unknown objects, we selected ten 3D models from the Princeton Shape Benchmark [34] and the Stanford 3D Scanning Repository [35] as known objects, and ten comparable, but distinct, other objects from the same databases as novel objects. They are all scaled to fit in a 6 $\times$ 6 $\times$ 6 bounding box. This should make it less likely for the framework to identify the correct object class merely by the length of the object. For this reason, the distance unit in this study does not correspond to real world measurements. An overview of the objects used in the experiments are shown in Fig. 4 and 5. Both known and novel objects are used as test objects in the experiment.
@@
For each trial in the experiment, a test object is initialized with a random pose and assumed to stay in that fixed position during the experiment. Once the first contact is established, the robotic arm will start to explore the object surface actively with the proposed exploration procedure, while updating it’s belief of the object class and object pose, and constantly reconstructing the shape until the termination criterion is satisfied. To explore the performance of the framework, ten trials are carried out per test objects. As performance indices we record object class prediction error, pose estimation error, reconstruction error, and termination time step.
@@ IV-B Results
For clarity, the analysis of the results are separated for known test objects and novel test objects. For known objects, pose error is measured by the TWD distance between the vertices of the MLE shape $\boldsymbol{M}$ and the vertices of the ground truth object shape $\boldsymbol{M}$. Similarly, for novel objects, the reconstruction error is measured by the TWD between the vertices of the GPIS $\boldsymbol{\boldsymbol{S}^{*}}$ and the vertices of the ground truth object shape $\boldsymbol{M}$. The surface uncertainty for both known and novel object is instead measured by the DHD between $\boldsymbol{M}$ and all contact points $\boldsymbol{X}_{c}$. An overview of the results for the experiments can be found in Fig. 6.
@@ IV-B1 Object recognition and pose estimation with known objects
When addressing known objects, regardless of the choice of exploration procedure, the framework correctly identified the class of the objects in all 100 trials and achieved pose estimation errors below the desired threshold of 0.6 in 100 out of 100 trials.
Figure 6: An overview of simulation experiment results, the first row is for known objects and the second row is for novel objects: (a) Average pose estimation error over time with GPIS-DHD and RRT-based exploration procedures for known objects. The orange and blue line indicate the mean pose error at each time step over 100 trials for the ten known objects. The red dashed line indicates the desired pose error threshold. (b) The time steps to reach below 0.6 pose error for each known object class. (c) Average DHD from MLE to contact data points over time with GPIS-DHD and RRT-based exploration procedures. The DHD from MLE shape to contact data points is used to determine if the contact data points cover the estimated surface sufficiently. The red dashed line indicates the DHD threshold for termination. (d) Termination time steps for known objects with GPIS-DHD and RRT-based exploration procedures. (e) Average shape reconstruction error for novel objects with GPIS-DHD and RRT-based exploration procedures for unknown objects. The red dashed line indicates the desired reconstruction error threshold. (f) The time steps to reach below 0.6 reconstruction error for each novel object class. (g) Average DHD from MLE to contact data points over time with GPIS-DHD and RRT-based exploration procedures for unknown objects. The DHD from GPIS to contact data points is used to determine if the contact data points cover the estimated surface sufficiently. The red dashed line indicates the DHD threshold for termination. (h) Termination time step for novel objects with GPIS-DHD and RRT-based exploration procedures. The translucent bands around the curves in (a) (c) (e) (g) and the error bars in (b) (d) (f) (h) indicate the confidence interval of 95%.
@@
Fig. 6b shows that the GPIS-DHD exploration had a small lead in reaching the desired pose estimation error in 9 out of 10 classes, while having a major lead in the class mug 1, compared to the RRT-based exploration procedure. With Fig. 7, one can see there is a clear difference of the pose estimation error over time between the two exploration procedures for the mug 1 class. To distinguish the pose of the mug, its handle must be located to break the rotational symmetry of its main body. The RRT-based exploration procedure on average took longer to locate the mug’s handle compared to the GPIS-DHD exploration procedure. The visualization of this special case can be found in Fig. 8.
@@
Figure 8: Comparison of the mug 1 case with GPIS-DHD and RRT exploration procedures respectively: On the left, GPIS-DHD exploration procedure managed to cover the mug’s handle after 200 contact points (red dots), therefore the ground truth (blue shape) and the MLE shape (red shape) overlapped. On the right, RRT exploration procedure failed to find the handle after 200 contact points, therefore lacked information to determine the mug’s pose.

### v3

@@ III-F Active Data Acquisition
As tactile data points at each time step only cover a fraction of the object surface, active data acquisition is required to address the uncertainty and ambiguity due to incomplete information. The process is divided into two steps: target point selection and contact enforcement. The idea is to first determine a target point, and then start exploration from the candidate target point until a contact point is found.
III-F1 Target Point Selection
@@
One way to select the target point is to calculate the posterior variance for each vertex on the reconstructed surface based on (30). Then the point with the highest posterior variance will be selected as the target point $\overline{\mathbf{x}}_{t+1}$ for the next time step, i.e.,
$\overline{\mathbf{x}}_{t+1}=\operatorname*{\text{arg max}}_{\mathbf{x}^{*}\in\mathbf{S}^{*}}\left\{\text{var}(\mathbf{x}^{*})\right\},$ (38)
$\text{var}(\mathbf{x}^{*})=\boldsymbol{\sigma}_{p}(\mathbf{x}^{*})_{0,0},$ (39)
$\mathbf{S}^{*}\subset\left\{\mathbf{x}^{*}\in R^{3}\;|\;\boldsymbol{\mu}_{p}(\mathbf{x}^{*})=0\right\},$ (40)
@@
with $\mathbf{S}^{*}$ a finite subset of the zero-level set of the posterior GPIS, which can be obtained through the marching cubes algorithm [31], and $\boldsymbol{\sigma}_{p}(\mathbf{x}^{*})_{0,0}$ denotes the first element of $\boldsymbol{\sigma}_{p}(\mathbf{x}^{*})$ in (30).
@@
However, this exploration procedure requires the framework to update the GPIS every time step, which is computationally expensive. To alleviate this issue, the GPIS will not be updated and the framework will use the MAP shape from the PF as estimated shape when the object is considered to be a known object. An exploration based on the directed Hausdorff distance (DHD) is used to guide the exploration when variance from the GPIS is not available. The point of interest for the next time step $\overline{\mathbf{x}}_{t+1}$ can be selected on the MAP surface by finding the point that yields the largest DHD, i.e.,
$\overline{\mathbf{x}}_{t+1}=\operatorname*{\text{arg max}}_{\mathbf{x}_{m}\in\mathbf{M}}\min_{\mathbf{x}_{c}\in\mathbf{X}_{c}}\lVert\mathbf{x}_{m}-\mathbf{x}_{c}\rVert_{2},$ (41)
@@
where $\mathbf{M}$ represents the set of all vertices of the MAP shape, and $\mathbf{X}_{c}$ denotes the set of observed contact points. The point of interest obtained through (41) is the point on the MAP shape with the largest distance to its nearest data point. And DHD is defined as follows,
$\mathbf{d}_{H}(\mathbf{A},\mathbf{B})=\max_{\mathbf{a}\in\mathbf{A}}\min_{\mathbf{b}\in\mathbf{B}}\lVert\mathbf{a}-\mathbf{b}\rVert_{2}$ (42)
where $\mathbf{A}$ and $\mathbf{B}$ are two point sets.
@@
III-F2 Contact Enforcement
With equations (38) and (41), a candidate target point is obtained at each time step, but there is no guarantee that it will lead to contact. On the contrary, since the target point is near the furthest point on the estimated surface from the existing data points, it often fails to establish contact at the proposed target location. A solution to guarantee contact at each time step is to perform contour following, namely, the tactile sensor could move towards the target point locally on the tangent plane at each time step. However, this approach might take a long time to reach the target location and may fail to capture the global context efficiently. As the selected target point is either on the GPIS or the MAP shape, confirming the contact near the target point can provide information that could greatly impact the belief of the framework. Negative information (non-contact points) in penetration with the estimated shape implies the current estimation is likely to be wrong, whereas positive information (contact) provides evidence for the current estimation.
Figure 3: Example of the contact enforcement procedure. Starting from the target point (red dot), the sensor first moves towards the interior of the MAP shape (purple dot). If no contact is found, a non contact point is recorded and the sensor moves towards the closest existing contact point (green dot). If no contact is established either, the closest known contact point is contacted again. From there, the algorithm takes small steps on the surface along the local tangent plane, towards the target point while remaining in contact with the surface. After a short distance, a new contact point will ultimately be recorded (yellow dot).
Fig. 3 sketches the contact enforcement procedure; see caption for details. During the contact enforcement procedure, most waypoints only contain negative information. Taking all the negative information data into account will lead to the negative information dominating the likelihood function, while the contact points would have less impact. To circumvent this, only one non-contact point can be registered at each time step. In this manner, the number of data points that contain negative information is no larger than the number of data points with positive information.
Alternatively, when no prior is available, an exploration procedure based on rapidly exploring random tree (RRT) can be adopted. It simply chooses a random point in the workspace, and it expands in that direction from the closest existing contact point. This requires neither prior knowledge nor global shape estimation as the expansion is local. The RRT-based exploration procedure is used as the fallback strategy when the GPIS-DHD exploration procedure fails to obtain new contact points.
III-G Termination Criterion
@@
$\mathbf{d}_{H}(\mathbf{M},\mathbf{X}_{c})\leq\epsilon$ (43)
The algorithm terminates if $\mathbf{d}_{H}(\mathbf{M},\mathbf{X}_{c})$ is lower than a small value $\epsilon$, where $\mathbf{M}$ represents the set of all vertices of the MAP shape, and $\mathbf{X}_{c}$ denotes the set of observed contact points. Intuitively, if any point on the estimated surface is at most $\epsilon$ away from its closest existing contact point, the program terminates. In other words, to terminate the exploration, the contact points have to cover the estimated object surface with a certain density related to $\epsilon$. The threshold $\epsilon$ can be interpreted as a level of detail parameter: lower $\epsilon$ results in more data being collected, albeit with longer exploration time. Identically, for novel objects, the termination criterion is
$\mathbf{d}_{H}(\mathbf{\mathbf{S}^{*}},\mathbf{X}_{c})\leq\epsilon,$ (44)
with $\mathbf{S}^{*}$ as the zero level set of the GPIS. Although the posterior variance of the GPIS can also be used for the same purpose, the proposed criterion $\mathbf{d}_{H}(\mathbf{\mathbf{S}^{*}},\mathbf{X}_{c})$ has the advantage of being independent of the choice of kernel parameters of the Gaussian process and the type of test objects at hand.
@@
  Initialization: Initialize the particle filter with the first contact point $\mathbf{x}_{1}$ and tactile observation $\mathbf{d}_{1}$.
  while $\mathbf{d}_{H}(\mathbf{\mathbf{S}^{*}},\mathbf{X}_{c})>\epsilon$ do
     Update belief using the particle filter, extract the MAP particle $(\mathbf{z}^{*},\overline{w}^{*})$ and its model evidence $p(\mathbf{D}|\mathbf{z}^{*},\mathbf{X})$
     if $p(\mathbf{D}|\mathbf{z}^{*},\mathbf{X})<(\frac{1}{Z}0.90)^{n_{pos}}*(0.50)^{n_{neg}}$ then
        Update GPIS and select the next target point $\overline{\mathbf{x}}_{t+1}$ from the zero level set of GPIS $\mathbf{S}^{*}$ using (38)
@@
        Select the next target point $\overline{\mathbf{x}}_{t+1}$ from the set of MAP surface $\mathbf{M}$ using(41)
@@
     Establish contact using the exploration procedure and add observed location to $\mathbf{X}$ and tactile observation to $\mathbf{D}$.
     Update $\mathbf{d}_{H}(\mathbf{\mathbf{S}^{*}},\mathbf{X}_{c})$ based on (42)
@@
III-H Shape Similarity Metric
@@
$\mathbf{d}_{TH}(\mathbf{A},\mathbf{B})=\max(\mathbf{d}_{H}(\mathbf{A},\mathbf{B}),\mathbf{d}_{H}(\mathbf{B},\mathbf{A}))$ (45)
The pose estimation error is measured by the TWD between the MAP shape from the PF and the actual shape of the test object. Similarly, the shape reconstruction error is measured by the TWD between the reconstructed shape from the GPIS and the actual shape of the test object.
@@ IV-A Experiment Setup
To show that the framework can address both known and novel objects, we selected ten 3D models from the Princeton Shape Benchmark [32] and the Stanford 3D Scanning Repository [33] as known objects, and ten comparable, but distinct, other objects from the same databases as novel objects. They are all scaled to fit in a 6 $\times$ 6 $\times$ 6 bounding box. This should make it less likely for the framework to identify the correct object class merely by the length of the object. For this reason, the distance unit in this study does not correspond to real world measurements. An overview of the objects used in the experiments are shown in Fig. 4 and 5. Both known and novel objects are used as test objects in the experiment.
@@
For each trial in the experiment, a test object is initialized with a random pose and assumed to stay in that fixed position during the experiment. At the beginning of each trial, the first contact point is sampled randomly on the test object’s surface. Once the first contact is established, the framework will start to explore the object surface actively with the proposed exploration procedure, while updating its belief of the object class and object pose, and constantly reconstructing the shape until the termination criterion is satisfied. To explore the performance of the framework, ten trials are carried out per test objects. As performance indices we record object class prediction error, pose estimation error, reconstruction error, and termination time step.
@@ IV-B Results
For clarity, the analysis of the results are reported separately for known test objects and novel test objects. For known objects, pose error is measured by the TWD between the MAP shape $\mathbf{M}$ and the ground truth object shape $\mathbf{G}$. Similarly, for novel objects, the reconstruction error is measured by the TWD between the reconstructed GPIS $\mathbf{S}^{*}$ and the ground truth object shape $\mathbf{G}$. The surface uncertainty for both known and novel object is instead measured by the DHD between $\mathbf{M}$ and all contact points $\mathbf{X}_{c}$. An overview of the results for the experiments can be found in Fig. 6.
@@ IV-B1 Object Recognition and Pose Estimation with Known Objects
When addressing known objects with the GPIS-DHD exploration procedure, the framework correctly identified the class of the objects in all 100 trials and achieved pose estimation errors below the desired threshold of 0.6 in 100 out of 100 trials. With the RRT-based exploration procedure, only in one case the framework failed to achieve the desired pose estimation error threshold, where no contact points were acquired on the handle of the mug.
Figure 6: An overview of simulation experiment results, the first row is for known objects and the second row is for novel objects: (a) Average pose estimation error over time with GPIS-DHD and RRT-based exploration procedures for known objects. The orange and blue line indicate the mean pose error at each time step over 100 trials for the ten known objects. The red dashed line indicates the desired pose error threshold. (b) The time steps to reach below 0.6 pose error for each known object class. (c) Average DHD from MAP to contact data points over time with GPIS-DHD and RRT-based exploration procedures. The DHD from MAP shape to contact data points is used to determine if the contact data points cover the estimated surface sufficiently. The red dashed line indicates the DHD threshold for termination. (d) Termination time steps for known objects with GPIS-DHD and RRT-based exploration procedures. (e) Average shape reconstruction error for novel objects with GPIS-DHD and RRT-based exploration procedures for novel objects. The red dashed line indicates the desired reconstruction error threshold. (f) The time steps to reach below 0.6 reconstruction error for each novel object class. (g) Average DHD from MAP to contact data points over time with GPIS-DHD and RRT-based exploration procedures for novel objects. The DHD from GPIS to contact data points is used to determine if the contact data points cover the estimated surface sufficiently. The red dashed line indicates the DHD threshold for termination. (h) Termination time step for novel objects with GPIS-DHD and RRT-based exploration procedures. The translucent bands around the curves in (a) (c) (e) (g) and the error bars in (b) (d) (f) (h) indicate the confidence interval of 95%. Despite the x axis in (b) (d) (f) (h) were clipped at 300 time steps for easier comparison, the experiments for known objects were stopped if they took beyond 300 time steps and the experiments for novel objects were stopped if they took beyond 400 time steps.
@@
Fig. 6b shows that the GPIS-DHD exploration had a small lead in reaching the desired pose estimation error in 9 out of 10 classes, while having a major lead in the class mug 1, compared to the RRT-based exploration procedure. In Fig. 7, a clear difference of the pose estimation error over time between the two exploration procedures can be seen for the mug. To distinguish the pose of the mug, its handle must be located to break the rotational symmetry of its main body. The RRT-based exploration procedure on average took longer to locate the mug’s handle compared to the GPIS-DHD exploration procedure. The visualization of this special case can be found in Fig. 8.
@@
Figure 8: Comparison of the mug 1 case with GPIS-DHD and RRT exploration procedures respectively: On the left, GPIS-DHD exploration procedure managed to cover the mug’s handle after 200 contact points (red dots), therefore the ground truth (blue shape) and the MAP shape (red shape) overlapped. On the right, RRT exploration procedure failed to find the handle after 200 contact points, therefore lacked information to determine the mug’s pose.